1. Solve the quadratic equation using the square root method: x^2=225
2. Solve the quadratic equation using the square root method: x^2-2=142
3. What are the dimensions of a square whose area is 36x^2

Answer by addingup(3677)

1. x^2 = 225 take the square root of both sides: x = 15 or x = -15
2. x^2-2=142; x^2 = 144 take the sqrt. both sides: x = 12 or x = -12
-------------------------------------
3. I don't understand what you're looking for. The sides of the square = the square root of the area. Is 36 the area? Then, what is x^2?